### v1

@@ Abstract
Availability and implementation: Simulation code for all models, data analyses, parameter inference and model selection is available online at https://github.com/rmcdomaths/tms/.
@@ 1 Introduction
Given multiple mathematical models which aim to reproduce the same biological data, determining which model and parameters give the best fit presents a theoretical and computational challenge. For example, spatio-temporal models often simulate complex high-dimensional data which is difficult to quantify and compare to observed data. Such models do not in general yield tractable likelihood functions, which significantly hinders the use of exact methods for parameter inference and model selection (Kirk et al., 2013).
Many mathematical models have been developed to study the mechanisms underlying tumour-induced angiogenesis (Scianna et al., 2013; Vilanova et al., 2017), a hallmark of cancer (Hanahan and Weinberg, 2011). Tumour cells use chemical signals to stimulate the growth of new blood vessels from existing vasculature (Ferrara, 2002), which provide a tumour mass with oxygen and nutrients that it requires to grow. However, instead of concise equations determining the growth of angiogenic networks, such models often comprise multiple agents and heterogeneous environments whose interactions depend non-deterministically on their spatial organisation. Discrete models of tumour-induced angiogenesis, for example, use multiple model rules and parameters to reproduce the branches, loops, and multiple components that characterise real vascular networks.
We use Topological Data Analysis (TDA), Approximate Bayesian Computation (ABC), and Random Forests (RFs) to develop a pipeline for parameter inference and model selection applicable to spatio-temporal models. TDA offers a toolkit of methods for quantifying spatial data and has previously been used to study multi-agent temporal systems (Topaz et al., 2015; Bhaskar et al., 2019; Stolz et al., 2024). In related work, TDA has been used to compare models of insect locomotion (Ulmer et al., 2019) and pattern formation in zebrafish (Cleveland et al., 2023). ABC provides a statistical framework for using model simulations to approximate posterior distributions when likelihood functions are not available (Lintusaari et al., 2016). RFs are an ensemble estimation method from machine learning which have previously been combined with ABC to estimate parameter values (Raynal et al., 2018) and rank candidate models (Pudlo et al., 2015).
@@ 2.1 Angiogenesis models
Discrete models of tumour-induced angiogenesis simulate the movement of individual Endothelial Cells (ECs). Vascular Endothelial Growth Factors (VEGF) produced by tumour cells initiate a cascade of chemical reactions which drive ECs towards the tumour. Early models employed the snail-trail model (Balding and McElwain, 1985) in which tip ECs migrate up spatial gradients of VEGF and fibronectin, while stalk ECs proliferate in their path to produce a contiguous line of cells. When two separate trails of ECs meet they fuse together to form a loop, in a process known as anastomosis. A tip EC may also split into two tip ECs, which thereafter move independently. Migrating, branching and looping tip ECs eventually reach the tumour mass and the connecting trails of stalk EC form a blood vessel network. Recent models reflect modern discoveries of cell mixing and phenotype switching (Stepanova et al., 2024), where ECs change type and overtake each other before forming a stable blood vessel network. Other models view ECs as a continuous population density rather than individual cells (Martinson et al., 2021) and account for blood flow and nutrient delivery when simulating vascular networks (Hormuth et al., 2021). We develop our pipeline of parameter inference and model selection on discrete angiogenesis models due to their simulation of finely resolved spatial data.
The Anderson-Chaplain (AC) (Anderson and Chaplain, 1998), Stokes-Lauffenberger (SL) (Stokes et al., 1991), and Plank-Sleeman (PS) (Plank and Sleeman, 2004) models employ the snail-trail model to simulate movement of individual ECs in a two-dimensional, square domain. We assume that VEGF levels increase from the bottom of the domain to a tumour at the top, guiding ECs to move upwards. Each model initalises multiple distinct tip ECs along the bottom of the domain, simulating their trajectories according to model-specific movement rules. Figure 1 illustrates the movement rules and model parameters in each model. See Supplementary Information Section 7 for full statements of each model and their parameters.
@@ 2.2 Data Generation and Analysis
Each angiogenesis model described in Section 2.1 simulates EC movement in a square domain, and we wish to quantify the spatial properties of each simulation. To do so, we overlay a regular grid onto the simulation domain at the final simulation timestep and compute a collection of statistics on the point cloud consisting of the grid locations of ECs. We compute the mean, standard deviation, minimum, maximum, range, interquartile range, and the 10th, 25th, 75th and 90th percentiles of the $x$ and $y$ co-ordinates of the EC grid locations. These spatially-averaged statistics measure how simulated ECs are distributed along the vertical and horizontal axes of the simulation domain, and were previously used for comparison to topological statistics by Nardini et al. (2021) to distinguish the AC model’s behaviour in different parameter regimes. To quantify spatial structures in simulated vasculature which may be obscured by spatially-averaged statistics (such as loops, branches and separated components) we use Extended Persistent Homology (EPH).
EPH arises from Persistent Homology (PH), a common pipeline within TDA for quantifying spatial features in data. To compute PH, spatial data must be converted into a sequence of simplicial complexes known as a filtration (see, for example, Otter et al. (2017)). Intuitively, a simplicial complex is a graph that includes nodes, edges, and higher-order connections such as triangles or tetrahedra. A filtration is a sequence of embedded simplicial complexes $\Sigma_{k}$ such that $\Sigma_{k}\subseteq\Sigma_{k+1}$ for $k=0,1,\dots,K-1$. The dimension-$p$ PH of a filtration is the sequence (1) of $\mathbb{F}$-vector spaces $H_{p}(\Sigma_{k})$ referred to as homology groups. Throughout this analysis, we work with the field $\mathbb{F}=\mathbb{Z}/2\mathbb{Z}$, which is widely adopted in applications for its simplicity and interpretability. The homology group $H_{p}(\Sigma_{k})$ captures topological features of dimension $p$ in $\Sigma_{k}$. Intuitively, $H_{0}$ detects connected components, $H_{1}$ detects loops and, in general, $H_{p}$ detects $p$-dimensional holes (Carlsson, 2009). The inclusion maps between successive simplicial complexes in the filtration induce linear maps in sequence (1) which allow us to track $p$-dimensional holes across the filtration.
$H_{p}(\Sigma_{0})\rightarrow\dots\rightarrow H_{p}(\Sigma_{k})\rightarrow\dots\rightarrow H_{p}(\Sigma_{K})$ (1)
Sequence (1) uniquely decomposes which allows us to represent each topological feature as a persistence pair $(b,d)$ (Zomorodian and Carlsson, 2005). The birth $b$ corresponds to the filtration index $k$ in the sequence at which a $p$-dimensional hole is first created in the simplicial complex, and the death $d$ is the filtration index in which this hole is filled in. The persistence $d-b$ measures how long the corresponding feature persists through the filtration. Note that persistence may be infinite if the $p$-dimensional hole is never filled in (and $d=\infty$). The interpretation of $(b,d)$ depends on how the filtration was constructed from the underlying spatial data. For example, Nardini et al. (2021) used a sweeping-plane filtration to analyse the AC model, in which persistence pairs quantified the location of loops and components in simulated vascular networks (measured in distance from the tumour).
Instead of using a standard PH pipeline, we compute Extended Persistent Homology (EPH) (Cohen-Steiner et al., 2009). Like PH, EPH uses a filtration to compute the sequence (1). In contrast to PH, EPH uses additional simplicial complexes to extend the sequence of vector spaces providing refined topological descriptors of spatial data. Specifically, EPH uses a sequence of simplicial complexes $\Sigma^{k}$ containing those simplices that are in $\Sigma_{K}$ but not in $\Sigma_{k}$. The relative homology $H_{p}(\Sigma_{K},\Sigma^{k})$ detects topological features in the quotient complex $\Sigma_{K}/\Sigma^{k}$ (Edelsbrunner and Harer, 2022), and since $\Sigma^{k}\supseteq\Sigma^{k+1}$ for $k=K-1,\dots,1,0$, quotient maps between these complexes induce linear maps tracking topological features through sequence (2).
$H_{p}(\Sigma_{K},\Sigma^{K})\rightarrow\dots\rightarrow H_{p}(\Sigma_{K},\Sigma^{k})\rightarrow\dots\rightarrow H_{p}(\Sigma_{K},\Sigma^{0})$ (2)
The concatenation of sequences (1) and (2) is known as the dimension-$p$ EPH of the filtration. $p$-dimensional holes may be detected and then filled within either sequence (1) or (2) and the corresponding indices $k$ may be extracted as persistence pairs $(b,d)$. We compute dimension $p=0$ and $p=1$ EPH of two sweeping-plane filtrations to quantify spatial features in data simulated by the AC, SL or PS models. Using these filtrations, extended persistence pairs quantify the size and location of components and loops in simulated vasculature, in addition to finding branches which emerge from vessel components when ECs bifurcate. Each sweeping-plane filtration measures these topological features in either the horizontal ($x$) or vertical ($y$) co-ordinate direction. See Supplementary Information Section 8 for details of the filtrations we use to compute EPH and an example computation on an angiogenesis simulation. We convert extended persistence pairs in each dimension into fixed-length vectors using two common vectorizations. Persistence Images (Adams et al., 2017) assign a weight to each extended persistence pair and create a smooth surface which is discretised over a grid to produce a stable, fixed-length vector. Persistence statistics (Ali et al., 2023) are vectors that consist of the mean, standard deviation, median, interquartile range, full range, and the 10th, 25th, 75th and 90th percentiles computed from the births $b$, deaths $d$, persistences $d-b$ and midpoints $(b+d)/2$ of persistence pairs $(b,d)$.
Figure 1: In the AC model, a tip EC makes one of five possible moves on a square lattice in each time-step according to probabilities $P_{0},P_{1},P_{2},P_{3},P_{4}$. A chemotaxis parameter biases movement probabilities in the direction of increasing VEGF concentration, and a haptotaxis parameter biases moves in the direction of increasing fibronectin. In the SL model, tip ECs move in any direction (off-lattice) with velocities modelled by a two-dimensional Stochastic Differential Equation (SDE). Parameters $\kappa$ and $\sigma$ determine how strongly an EC’s current velocity $w$ is affected by the VEGF gradient $c$ and random variation $r$ respectively. The PS model assigns a constant speed to each EC and, at each time-step, rotates the angle that the velocity vector makes with the vertical by $\hat{\phi}$. The probability $\hat{\tau}_{n}^{+}+\hat{\tau}_{n}^{-}$ of turning is determined by a turning rate parameter, and a chemotaxis parameter biases turns that re-orient the EC’s direction towards the tumour. In all models, a tip EC may bifurcate into two ECs which thereafter move independently if its age exceeds the minimum age for branching parameter $a_{\text{br}}$ and the VEGF concentration at its location exceeds the the VEGF threshold for branching parameter $c_{\text{br}}$. We show how many spatially-averaged and topological statistics, computed in either the $x$ or $y$ co-ordinate direction, appear among the $100$ most important summary statistics to the inference of each parameter.
@@ 3.1 Approximate Bayesian Computation
where the evidence $p(\mathcal{D})$ is the integral $\int_{\Theta}p(\mathcal{D}|\Theta)p(\Theta)\textnormal{d}\Theta$ over parameters $\Theta$ in the support of the prior. Although Bayes’ rule gives a closed formula for the posterior distribution, it is often impractical to use directly. The likelihood function $p(\mathcal{D}|\Theta)$ may be too complicated to derive for probabilistic spatial models in which many datasets $\mathcal{D}$ may be simulated from the same parameters $\Theta$. Instead, Bayes’ rule is used to derive Algorithm 1 which allows sampling from the posterior, even when the likelihood and evidence are not known. When the tolerance $\epsilon$ in Algorithm 1 is set to zero, the distribution of accepted parameters is (a constant multiple of) the posterior distribution $p(\Theta|\mathcal{D}^{*})$ (Frazier et al., 2018). However, it is often not appropriate or possible to seek an exact posterior distribution from observed data, since a model may rarely reproduce observed data $\mathcal{D^{*}}$ exactly, and the observed data may be noisy. It is therefore advisable to choose $\nu$ and $\epsilon$ such that parameter values are accepted if they simulate data that is similar to observed data. The general form of such a distance function is $\nu(\mathcal{D}^{*},\mathcal{D}_{i})=\norm{X^{*}-X_{i}}_{2}$ where $X_{i}$ and $X^{*}$ are vectors of summary statistics computed from model data $\mathcal{D}_{i}$ and observed data $\mathcal{D}^{*}$ respectively (this distance may be averaged over several instances of observed data $\mathcal{D}^{*}$). Summary statistics aim to capture relevant properties of data as a low dimensional vector. As the tolerance $\epsilon$ approaches 00, the distribution of parameters accepted by Algorithm 1 approaches $p(\Theta|X^{*})$, which equals $p(\Theta|\mathcal{D}^{*})$ if the summary statistics are sufficient for the model in question, or is a close approximation if the summary statistics are insufficient but informative (Joyce and Marjoram, 2008). However, if $\epsilon$ is too small, few parameters will be accepted and many simulations may be needed to approach the true posterior. To overcome the difficulty of choosing a single tolerance, ABC with Sequential Monte Carlo sampling (ABC-SMC) modifies Algorithm 1, and instead generates a sequence of immediate distributions which approach the true posterior using tolerances $\epsilon_{0}>\epsilon_{1}>\dots>\epsilon_{\textnormal{final}}>0$. For example, the ABC-SMC algorithm of Del Moral et al. (2012) generates the schedule of tolerances by decreasing the proportion of accepted parameters in each intermediate distribution by a fixed factor $\alpha\in(0,1)$. We discuss the choice of summary statistics and $\nu$ in Section 4.
Algorithm 1 ABC with rejection sampling
3:for Candidate parameter values $\theta_{i}$ sampled from $p(\Theta)$ do
5:   if $\nu(\mathcal{D}^{*},\mathcal{D}_{i})\leq\epsilon$ then
6:      accept $\theta_{i}$
7:   end if
8:end for

### v2

@@ Abstract
Availability and implementation: Simulation code for all models, data analyses, parameter inference and model selection is available online at https://github.com/rmcdomaths/tms/ and archived at https://doi.org/10.5281/zenodo.17392787.
@@ 1 Introduction
Given multiple mathematical models which aim to reproduce the same biological data, determining which model and parameters give the best fit presents a theoretical and computational challenge. For example, spatio-temporal models often simulate complex high-dimensional data which is difficult to quantify and compare to observed data. Such models do not in general yield tractable likelihood functions, which significantly hinders the use of exact methods for parameter inference and model selection (modelSelectionReview).
Many mathematical models have been developed to study the mechanisms underlying tumour-induced angiogenesis (preziosireview; vilanovareview), a hallmark of cancer (hallmarks). Tumour cells use chemical signals to stimulate the growth of new blood vessels from existing vasculature (VEGFforAngio), which provide a tumour mass with oxygen and nutrients that it requires to grow. However, instead of concise equations determining the growth of angiogenic networks, such models often comprise multiple agents and heterogeneous environments whose interactions depend non-deterministically on their spatial organisation. Discrete models of tumour-induced angiogenesis, for example, use multiple model rules and parameters to reproduce the branches, loops, and multiple components that characterise real vascular networks.
We use Topological Data Analysis (TDA), Approximate Bayesian Computation (ABC), and Random Forests (RFs) to develop a pipeline for parameter inference and model selection applicable to spatio-temporal models. TDA offers a toolkit of methods for quantifying spatial data (ghristBarcodes; carlssonTDA; edelsbrunnerHarer). TDA has previously been used to study multi-agent temporal systems (tdaBiologicalAggregation; TDAcollectivemotion; relationalPH) and was used in related work to compare models of insect locomotion (TDAbiomodels) and pattern formation in zebrafish (sandstedeModelTDA). ABC provides a statistical framework for using model simulations to approximate posterior distributions when likelihood functions are not available (ABCreview). RFs are an ensemble estimation method from machine learning which have previously been combined with ABC to estimate parameter values (ABCRF_paraminf) and rank candidate models (ABCRF_modelchoice).
@@ 2.1 Angiogenesis models
Discrete models of tumour-induced angiogenesis simulate the movement of individual Endothelial Cells (ECs). Vascular Endothelial Growth Factors (VEGF) produced by tumour cells initiate a cascade of chemical reactions which drive ECs towards the tumour. Early models employed the snail-trail model (originalSnailTrail) in which tip ECs migrate up spatial gradients of VEGF and fibronectin, while stalk ECs proliferate in their path to produce a contiguous line of cells. When two separate trails of ECs meet they fuse together to form a loop, in a process known as anastomosis. A tip EC may also split into two tip ECs, which thereafter move independently. Migrating, branching and looping tip ECs eventually reach the tumour mass and the connecting trails of stalk ECs form a blood vessel network. Recent models reflect modern discoveries of cell mixing and phenotype switching (angioCellMixing), where ECs change type and overtake each other before forming a stable blood vessel network. Other models view ECs as a continuous population density rather than individual cells (continuumangiomodels) and account for blood flow and nutrient delivery when simulating vascular networks (flowangiomodels). We develop our pipeline of parameter inference and model selection on discrete angiogenesis models due to their simulation of finely resolved spatial data.
The Anderson-Chaplain (AC) (andersonChaplain), Stokes-Lauffenberger (SL) (stokesLauffenberger), and Plank-Sleeman (PS) (plank_sleeman) models employ the snail-trail model to simulate movement of individual ECs in a two-dimensional, square domain. We assume that VEGF levels increase from the bottom of the domain to a tumour at the top, guiding ECs to move upwards. Each model initalises multiple distinct tip ECs along the bottom of the domain, simulating their trajectories according to model-specific movement rules. In each model, we choose four parameters that are likely to lead to measurable changes in simulated data and we attempt to infer their values. Figure 1 illustrates the movement rules and model parameters in each model. See Supplementary Information section 7 for full statements of each model and its parameters.
Figure 1: In the AC model (andersonChaplain), a tip endothelial cell (EC) makes one of five possible moves on a square lattice in each time-step according to probabilities $P_{0},P_{1},P_{2},P_{3},P_{4}$. A chemotaxis parameter $\chi$ biases movement probabilities in the direction of increasing VEGF concentration, and a haptotaxis parameter $\rho$ biases moves in the direction of increasing fibronectin. In the SL model (stokesLauffenberger) tip ECs move in any direction (off-lattice) with velocities modelled by a two-dimensional stochastic differential equation. Parameters $\kappa$ and $\sigma$ determine how strongly an EC’s current velocity $w$ is affected by the VEGF gradient $c$, and random variation $r$ respectively. The PS model (plank_sleeman) assigns a constant speed to each tip EC and, at each time-step, rotates the angle that the velocity vector makes with the vertical. The probability $\hat{\tau}_{n}^{+}+\hat{\tau}_{n}^{-}$ that a tip EC turns by $\hat{\phi}$ is determined by a turning rate parameter $D_{r}$. A chemotaxis parameter $d_{c}$ biases turns that re-orient the EC’s direction towards the tumour. In all models, a tip EC may bifurcate into two ECs which thereafter move independently if its age exceeds the minimum age for branching parameter $a_{\text{br}}$ and the VEGF concentration at its location exceeds the the VEGF threshold for branching parameter $c_{\text{br}}$. We show how many spatially-averaged and topological summary statistics, computed in either the $x$ or $y$ co-ordinate direction, appear among the $100$ most important summary statistics to the inference of each parameter.
@@ 2.2 Data Generation and Analysis
Each angiogenesis model outlined in section 2.1 (and described fully in Supplementary Information section 7) simulates EC movement in a square domain. To summarise the spatial properties of each simulation, we overlay a regular grid onto the domain at the final timestep and compute a collection of spatially-averaged and topological summary statistics. For the spatially-averaged summary statistics, we compute the mean, standard deviation, minimum, maximum, range, and the 10th, 25th, 75th and 90th percentiles of the $x$ and $y$ co-ordinates of EC locations in the grid. These were used in nardiniclustering to distinguish the AC model’s behaviour in different parameter regimes.
Persistent homology (PH) is a prominent method within Topological Data Analysis (TDA) (ghristBarcodes; carlssonTDA; edelsbrunnerHarer) to quantify loops, branches, and connected components. Here, we require finer information than is provided by standard persistence; therefore, we use extended persistent homology (EPH) (extendedpersistence). We give an overview of PH and EPH, briefly indicating how EPH arises from PH and captures a greater range of spatial information within the angiogenesis data we simulate. See Supplementary Information section 8 for a full definition of EPH and a worked example.
Figure 2: A persistence diagram (PD) and extended persistence diagram (EPD) for a simple blood vessel computed using the vertical sweeping-plane filtration. The PD points quantify the size and location of the small lower branch (
$\blacksquare$
), and the locations of the component (
$\blacksquare$
) and loop (
$\bullet$
). The EPD points quantify the location and size of all topological features quantified by the PD, in addition to the small upper branch (
$\bullet$
). The branches (
) are not detected by PH or EPH with this sweeping-plane filtration.
To compute PH, a nested sequence of simplicial complexes, known as a filtration, is built on the data. Intuitively, a simplicial complex $\Sigma_{k}$ is a graph that includes nodes and edges as well as higher-order connections such as triangles or tetrahedra. A filtration must be carefully constructed such that each $\Sigma_{k}$ encodes the spatial properties of the underlying data at some spatial threshold defined by $k$. For example, nardiniclustering used a sweeping-plane filtration to analyse the AC model, where $\Sigma_{k}$ is constructed from those ECs which are a distance of $k$ or more away from the tumour. Once a filtration has been chosen, one computes a sequence of $\mathbb{F}$-vector spaces $H_{p}(\Sigma_{k})$ known as homology groups. We use the field $\mathbb{F}=\mathbb{Z}/2\mathbb{Z}$, which is widely adopted in applications for its simplicity and interpretability. Homology groups quantify $p$-dimensional topological features in each $\Sigma_{k}$ (see, for example, roadmap). $H_{0}$ detects connected components, $H_{1}$ detects loops and, in general, $H_{p}$ detects $p$-dimensional voids. We consider dimensions 0 and $1$ only, since voids of dimension $2$ or higher do not appear in the angiogenesis data we simulate. Persistence pairs $(b,d)$ are computed from the sequence of homology groups to quantify topological features in the filtration (structurePH). A birth $b$ corresponds to the index $k$ in the filtration at which a topological feature first appears. A death $d$ is either the index $k$ at which the $p$-dimensional void it represents is filled in, or $\infty$ if the topological feature persists through the filtration. The difference dbd-b is known as the persistence of a topological feature. The interpretation of $(b,d)$ in terms of the underlying data depends on the choice of filtration. In the sweeping-plane filtration used in nardiniclustering, persistence pairs quantify connected components and loops in simulated vascular networks in terms of their distance from the tumour.
In PH, some topological features typically persist throughout the entire filtration. In Figure 2, for example, a connected component appears at $y=0$ and a loop at $y=0.25$, and both persist for all values of $y$ in the sweeping-plane filtration. The corresponding persistence pairs are therefore $(0,\infty)$ and $(0.25,\infty)$, which quantify limited location information and no size information about the topological features they represent. Furthermore, PH computes one PD in each dimension ($p=0$ and $p=1$ in this work), but we want to distinguish several different spatial structures (see Figure 2 for examples).
We would like to quantify the size and location of different spatial features. EPH provides this information by appending relative homology groups (edelsbrunnerHarer) to the sequence of ordinary homology groups in PH. Topological features which persist through all ordinary homology groups will die in the relative homology groups, so persistence pairs computed from EPH are guaranteed to have finite persistence. Each EPH persistence pair then be classified as one of four types depending on where the birth and death appear in the sequence of ordinary and relative homology groups, which provides additional information about the corresponding topological features. See Supplementary Information section 8 for a formal definition of topological feature types in EPH and their interpretation in simulated angiogenesis data. Figure 2 compares persistence pairs computed from a simple blood vessel network using PH and EPH, illustrating the extra information provided by EPH. We compute two extended persistence diagrams (EPDs) for each angiogenesis dataset–using a vertical $(y)$ and a horizontal $(x)$ sweeping-plane filtration. We vectorise each EPD using Persistence Images (persistenceimage) and persistence statistics (vectorisationSurvey), and our topological summary statistics are the concatenation of these vectors.
@@ 3.1 Approximate Bayesian Computation
where the evidence $p(\mathcal{D})$ is the integral $\int_{\Theta}p(\mathcal{D}|\Theta)p(\Theta)\textnormal{d}\Theta$ over parameters $\Theta$ in the support of the prior. Although Bayes’ rule gives a closed formula for the posterior distribution, it is often impractical to use directly. The likelihood function $p(\mathcal{D}|\Theta)$ may be too complicated to derive for probabilistic spatial models in which many datasets $\mathcal{D}$ may be simulated from the same parameters $\Theta$. Instead, Bayes’ rule is used to derive Approximate Bayesian Computation (ABC) algorithms (ABCreview) which allow sampling from the posterior when the likelihood and evidence are not known. ABC algorithms sample candidate parameters $\theta_{i}$ from the prior $p(\theta)$ and accept them if the distance $\nu(\mathcal{D}_{i},\mathcal{D}^{*})$ between simulated and observed data is less than some tolerance $\epsilon>0$ for some distance function $\nu$. If the tolerance $\epsilon$ is set to zero, then the distribution of accepted parameters is the posterior $p(\Theta|\mathcal{D}^{*})$ (ABCproof). However, it is often not appropriate or possible to seek an exact posterior distribution from observed data, since a model may rarely reproduce observed data $\mathcal{D^{*}}$ exactly, and the observed data may be noisy. It is therefore advisable to choose $\nu$ and $\epsilon$ such that parameter values are accepted if they simulate data that is similar to observed data. The general form of such a distance function is $\nu(\mathcal{D}^{*},\mathcal{D}_{i})=\norm{X^{*}-X_{i}}_{2}$ where $X_{i}$ and $X^{*}$ are vectors of summary statistics computed from model data $\mathcal{D}_{i}$ and observed data $\mathcal{D}^{*}$ respectively. Summary statistics aim to capture relevant properties of data as a low dimensional vector. As $\epsilon$ approaches 0, the distribution of parameters accepted by ABC algorithms approaches $p(\Theta|X^{*})$, which equals $p(\Theta|\mathcal{D}^{*})$ if the summary statistics are sufficient for the model in question, or is a close approximation if the summary statistics are insufficient but informative (sufficientStats).